# It’s High Time : A Survey of Temporal Information Retrieval and Question Answering

### v2

@@ 5 Approaches in Temporal IR/QA
A wide range of approaches have been developed to address the challenges of Temporal IR/QA, from early rule-based systems and statistical models to neural networks and large language models (LLMs). They differ in how they represent temporal information, reason over temporal relationships, and adapt to changing world knowledge.
5.1 Rule-based & Statistical Methods
Early work in Temporal IR/QA was dominated by rule-based systems and statistical models that laid the groundwork for core temporal tasks such as time expression normalization, event ordering, and temporal ranking. While limited in scalability and adaptability, they introduced many foundational concepts that remain relevant today.
Despite their simplicity, rule-based and statistical methods introduced key mechanisms of temporal intent modeling, expression normalization, and timeline reasoning that continue to influence more advanced systems.
@@ 5.2 Temporal Language Models
The emergence of deep learning has significantly advanced Temporal IR/QA by enabling models to capture temporal dependencies and contextual nuances. Recent research has led to the development of Temporal Language Models (TLMs) that explicitly incorporate temporal signals during pretraining or fine-tuning. Models such as TempoT5 Dhingra et al. (2022), TempoBERT Rosin et al. (2022), and BiTimeBERT Wang et al. (2023) included timestamps and temporal expressions directly into their training inputs or used time-focused pretraining tasks, improving temporal generalization in downstream tasks such as semantic change detection and Temporal QA. Other approaches, like syntax-guided temporal language model (SG-TLM) Su et al. (2023), enhance sensitivity to temporal structure by masking syntactic and semantic spans that carry temporal meaning.
On the other hand, Cao and Wang (2022) explored time-aware generation by introducing temporal prompts, including both natural language timestamp descriptions and continuous vector (linear) representations of timestamps. Beyond input-level integration, time-aware language models like TALM Ren et al. (2023) incorporate time-specific word representations through hierarchical modeling and temporal adaptation, achieving strong results in historical text dating. TCQA Son and Oh (2023) employs synthetic data and a time-context span selection task to train models that align time-aware representations with contextually grounded answers. Further, techniques such as Temporal Span Masking (TSM) Cole et al. (2023) and temporal attention mechanisms Rosin and Radinsky (2022) incorporate explicit temporal annotations into transformer architectures to improve time sensitivity.
5.2.1 Temporal RAG
@@
Recent temporal RAG systems extend this idea by embedding temporal signals directly into retrieval and generation pipelines. TempRALM Gade and Jetcheva (2024) introduces temporal signals into dense retrieval, enhancing recency and factual grounding for time-sensitive queries. TempRetriever Abdallah et al. (2025) and TsContriever Wu et al. (2024) encode temporal relevance directly into dense retrievers, improving alignment between temporal queries and evidence. TimeR4 Qian et al. (2024) proposes a Retrieve-Rewrite-Retrieve-Rerank pipeline that transforms implicit temporal queries into explicit ones, retrieves from time-anchored knowledge sources, and reranks based on temporal constraints. MRAG Siyue et al. (2024) adapts RAG with multi-source and multi-hop temporal retrieval for event-centric QA. To mitigate hallucinations and outdated generations, FRESHPROMPT Vu et al. (2024) integrates real-time signals into the prompting and retrieval process. Together, these models make RAG more responsive to temporal dynamics in IR/QA.
5.2.2 Temporal Reasoning Capabilities
While Temporal Language Models enhance time-aware representations and retrieval, many Temporal IR/QA tasks demand more sophisticated reasoning, such as understanding event sequences, temporal constraints, and durations.
Temporal reasoning capabilities in pre-trained language models (PLMs) have seen notable improvements, with recent efforts focusing on enhancing zero-shot generalization and temporal robustness. Continual temporal adaptation methods, including ECONET Han et al. (2021), enhance temporal relational coherence and consistency across evolving contexts. Structural temporal reasoning models like TIMERS Mathur et al. (2021), and ConTempo Niu et al. (2024) address multi-hop and document-level inference with specialized architectures. Moreover, event duration and ordering prediction have benefited from task-specific temporal pretraining objectives (e.g., E-PRED, R-PRED) Yang et al. (2020) and transfer learning strategies Virgo et al. (2022).
Despite these advancements, modeling temporal relationships in LLMs remains challenging. Recent benchmarks such as TRAM Wang and Zhao (2024) evaluate LLMs on tasks like event ordering, arithmetic, frequency, and duration, revealing that even strong models like GPT-4 fall short of human-level performance. To isolate genuine reasoning abilities from memorization, Test of Time (ToT) Fatemi et al. (2024) introduces synthetic tasks targeting temporal logic and inference. Additionally, TODAY Feng et al. (2023) challenges models with subtle temporal shifts and differential analysis. Methods like Narrative-of-Thought Zhang et al. (2024) guide models to generate structured temporal narratives. Finally, Wallat et al. (2024, 2025) study temporal blind spots of LLMs and their resiliency to changes in time-related elements (e.g., altering a date in a query, or its position) elucidating missing knowledge and showing that current models are still vulnerable to adversarial or other perturbations.
6 Future Directions
Despite two decades of research and significant progress, Temporal IR/QA systems still struggle to adapt to evolving real-world events, shifting user needs, and dynamic data streams. To advance the development of time-aware systems, we propose future directions organized into three core themes: System Design (architectures and real-time capabilities), Knowledge Management (updating and representing time-sensitive knowledge), and Evaluation and Robustness (metrics and generalization). These directions address gaps identified throughout this survey, such as temporal bias, rigidity in models, and the limited scope of existing evaluations.
6.1 System Design
Real-Time Information Integration. Most IR systems depend on periodically updated corpora, leaving them blind to rapidly unfolding events like elections, protests, or trending information. Future work should treat data as a continuous stream, enabling real-time indexing Baeza-Yates and Ribeiro-Neto (2011), burst detection Wang et al. (2021a), and responsive re-ranking Tran et al. (2015), as well as supporting applications like live event tracking or misinformation detection V et al. (2024).
Development of Temporally-Aware LLM Agents.
Current LLM agents prioritize task completion or dialogue but lack structured temporal understanding Wallat et al. (2024). Future systems should include dedicated temporal understanding methods for better understanding temporal references, semantics, and test-time reasoning.
6.2 Knowledge Management
Advanced Temporal Knowledge Editing. Static models struggle to keep up with real-world change. Instead of retraining, future systems could use modular, trackable edit layers for local updates, preserving historical facts.
Integration of Diachronic and Synchronic Knowledge.
Temporal questions often require combining evolving facts (e.g., event timelines) with stable knowledge (e.g., definitions). Future systems should integrate diachronic sources with synchronous sources to provide comprehensive answers. For example, answering "How has the unemployment rate changed since 2008?" requires diachronic trends from datasets like ArchivalQA (Section 3) and synchronous explanations from Wikipedia, addressing the aggregation needs (Section 2).
Multilingual Temporal IR/QA.
Temporal expressions vary across languages and cultures, posing challenges for globalized systems. For instance, date formats differ (e.g., DD/MM/YYYY vs. MM/DD/YYYY), and cultural references (e.g., "post-Meiji era" in Japanese) require context-specific interpretation. Future research should develop cross-lingual temporal taggers, multilingual benchmarks, and culturally adaptive models, building on multilingual taggers like HeidelTime Strötgen and Gertz (2010).
6.3 Evaluation and Robustness
Implicit Temporal Intent Understanding. Many queries imply but do not state a time frame. Future work should improve models’ ability to infer latent temporal scopes using derived labels or event grounding. This addresses the implicit reasoning challenges in datasets like TORQUE Ning et al. (2020) and TiQ Jia et al. (2024).
Robustness to Temporal Drift and Misalignment.
Performance drops when models are applied to data from different time periods, which can reduce accuracy Shin et al. (2025); Zhang and Choi (2023); Luu et al. (2022); Wallat et al. (2025). Future work should enhance model resilience to temporal misalignment, building on the robustness challenges in Test of Time.
@@ 7 Conclusion
Temporal IR/QA is critical for retrieving and reasoning over time-sensitive information in dynamic, evolving contexts. In this survey, we have traced the field’s progression from early rule-based systems to TLMs and RAG approaches. We identified core challenges, including temporal tagging, temporal intent detection, event ordering, and robustness to evolving facts and implicit temporal signals.
Our review highlights persistent limitations such as reliance on static knowledge, limited capabilities for future-oriented reasoning, and dataset bias toward past events. We show that temporal complexity, vague expressions, knowledge drift, and real-time demands significantly impact system behavior and evaluation.
Despite notable progress, current systems often struggle with temporal uncertainty, maintaining consistency across time, and adapting to multilingual or culturally diverse temporal expressions. As real-world applications increasingly require temporally adaptive systems, these gaps point to the need for richer evaluation protocols, improved temporal representations, and continual learning strategies. We anticipate future progress toward robust, time-aware IR and QA systems capable of understanding not just what happened, but also when, why, and how information evolves over time.
@@
This survey aims to provide a comprehensive overview of Temporal IR/QA. There are a few important limitations to acknowledge.
We made our best efforts to be thorough, but it is possible that some relevant works may have been missed. We conducted an extensive literature review using forward and backward snowballing techniques, with particular attention to papers published in major venues such as ACL, SIGIR, EMNLP, NeurIPS, ECIR, and preprints on arXiv. On the other hand, due to page limitations, we provide only a very brief summary of each method without exhaustive technical details.
Appendix A Appendix
A.1 Related Surveys
@@
While IR and QA have been widely surveyed, most existing reviews focus on general techniques, often neglecting temporal aspects. IR surveys emphasize ranking functions, neural retrieval models, and query understanding Li et al. (2025); Zhu et al. (2023), while QA surveys center on extractive, abstractive, or multi-hop answering over static knowledge sources Zhu et al. (2021); Mavi et al. (2024). They rarely consider temporal intent, dynamic or evolving information, or event sequencing, highlighting a key gap that remains unaddressed.
Several earlier works provided foundational insights into Temporal IR. Alonso et al. (2011) discusses challenges such as real-time streams, exploratory temporal search, and spatio-temporal retrieval. Campos et al. (2014) offers a broad overview of document dating, time-aware ranking, and query understanding, covering both explicit and implicit time signals. Kanhabua and Anand (2016) complements these with a tutorial on temporal indexing and ranking, emphasizing the detection of temporal query intent. There has been however no recent systematic overview of the field despite much research interest.
As a parallel research line, TQA over knowledge graphs has gained considerable attention Jia et al. (2021); Saxena et al. (2021); Chen et al. (2023); Xiong et al. (2024). Our survey focuses on temporally aware IR/QA over text. We review both traditional and neural approaches to core tasks such as temporal tagging, event dating, time-aware retrieval, and temporal reasoning. To our knowledge, no prior survey brings together recent developments across these tasks in the context of text-based IR/QA. Other related topics, including temporal fact verification Barik et al. (2024) and timeline summarization Sojitra et al. (2024), are discussed only when directly relevant.
@@
Temporal IR/QA, draw=black, fill=gray!10, thick, text=black, ver
@@
TIR Dataset ( §3.3),
color=orange!90, fill=orange!20, thick, text=black
Joho et al. (2014), Joho et al. (2016),
,color=orange!90, fill=orange!20, thick, text=black, leaf,text width=20em
Temporal Tasks ( §4),
color=cyan!100, fill=cyan!15, thick, text=black
Prediction,
@@
Das et al. (2017), Wang et al. (2021b)
Morbidoni et al. (2018),
Wang et al. (2021b)
, color=cyan!100, fill=cyan!15, thick, text=black, leaf
@@
Dalli (2006), Jatowt et al. (2007)
Kumar et al. (2012)
,, color=cyan!100, fill=cyan!15, thick, text=black, leaf
@@
Jatowt et al. (2013),
Shrivastava et al. (2017)
,, color=cyan!100, fill=cyan!15, thick, text=black, leaf
@@
Kanhabua and Nørvåg (2010),
Dakka et al. (2008),
Jones and Diaz (2007)
, color=cyan!100, fill=cyan!15, thick, text=black, leaf

### v3

@@ 4 Approaches in Temporal QA
A wide range of approaches have been developed to address the challenges of TQA, from early rule-based systems and statistical models to neural networks and large language models (LLMs). Tables 2 and 3 (see Appendix) provide a comprehensive comparison of these temporal QA approaches across architectural paradigms, temporal representations, and methodological strategies. They differ in how they represent temporal information, reason over temporal relationships, and adapt to changing world knowledge.
Effective TQA often relies on several core Temporal Prediction Tasks that support deeper temporal understanding. These tasks focus on inferring implicit or missing temporal signals from text, which is especially important when explicit temporal metadata is sparse or unreliable. Key tasks include Event Dating, Document Dating, Focus Time Estimation, Query Time Profiling, and Event Occurrence Prediction. They enable a range of downstream applications such as timeline construction, historical information retrieval, and time-sensitive question answering.
In this section, we focus on neural approaches, particularly temporal language models and retrieval-augmented generation (RAG) for TQA. For a more detailed discussion of temporal prediction tasks, as well as traditional rule-based and statistical QA methods, we refer the reader to Appendix C.
@@ 4.1 Temporal Language Models
The emergence of deep learning has significantly advanced TQA by enabling models to capture temporal dependencies and contextual nuances. Recent research has led to the development of Temporal Language Models (TLMs) that explicitly incorporate temporal signals during pretraining or fine-tuning. Models such as TempoT5 Dhingra et al. (2022), TempoBERT Rosin et al. (2022), and BiTimeBERT Wang et al. (2023) included timestamps and temporal expressions directly into their training inputs or used time-focused pretraining tasks. These strategies improve temporal generalization in downstream tasks such as semantic change detection and TQA.
Beyond timestamp inputs, more recent work introduces architectural modifications to capture temporal structure internally. TALM Ren et al. (2023) models time hierarchically, learning both short-term and long-term temporal word representations. SG-TLM Su et al. (2023) enhances temporal sensitivity by masking syntactic spans associated with time expressions, encouraging the model to learn implicit temporal cues. Similarly, models like Temporal Span Masking (TSM) Cole et al. (2023) and Temporal Attention Rosin and Radinsky (2022) explicitly incorporate temporal supervision into attention heads, enabling more fine-grained alignment between time expressions and relevant content.
Some TLMs focus on generation tasks that require temporal grounding. For instance, Cao and Wang (2022) proposes temporal prompts, both textual and continuous vector-based, that guide generation with time-specific context. TCQA Son and Oh (2023) introduces a synthetic QA dataset and a span-selection task that aligns answers with their temporal context, simulating historical reasoning and timeline consistency.
A common trend across TLMs is the shift from treating time as an auxiliary signal to making it a core component of the modeling process. Timestamp conditioning, temporal pretraining tasks and attention adaptation have improved temporal reasoning, but most models still rely on static corpora and struggle with vague or implicit expressions. TLMs represent a significant step forward but require further development to handle real-world complexity and temporal drift.
4.2 Temporal RAG
@@
A key trend in this space is the integration of temporal signals directly into the retrieval process. Models such as TempRetriever Abdallah et al. (2025) and TsContriever Wu et al. (2024) extend dense retrievers to account for temporal relevance by encoding queries and passages with timestamp-aware embeddings. These approaches improve alignment between temporally scoped questions and contextually valid evidence. Similarly, TempRALM Gade and Jetcheva (2024) enhances dense retrieval with temporal constraints to reduce factual drift and improve recency-grounded responses.
Other systems modify the retrieval generation pipeline itself. TimeR4 Qian et al. (2024) proposes a four-stage Retrieve-Rewrite-Retrieve-Rerank architecture that transforms underspecified temporal queries into explicitly time-anchored formulations before retrieving and re-ranking documents based on temporal fit. MRAG Siyue et al. (2024) addresses multi-hop reasoning across events by retrieving from multiple time-scoped sources and aggregating cross-temporal evidence for complexQA.
These models illustrate an emerging paradigm shift: from static temporal representation to real-time, adaptive temporal reasoning. Temporal RAG systems help mitigate hallucinations and better reflect recent developments, addressing temporal incompleteness. However, their success is tied to the quality and temporal alignment of the retriever, and they raise new challenges around evidence reliability and temporal grounding.
4.3 Temporal Reasoning Capabilities
While Temporal Language Models improve time-aware representation and retrieval, many TQA tasks require explicit temporal reasoning, such as tracking event sequences, resolving durations, handling temporal arithmetic, and drawing logical inferences across time. These capabilities are essential for questions involving change over time, causality, or evolving narratives, where retrieval alone is insufficient. .
Recent efforts have aimed to augment the reasoning capacity of pretrained language models (PLMs) through both architectural design and specialized training objectives. For example, ECONET Han et al. (2021) focuses on continual adaptation to changing temporal contexts, preserving consistency across updates. Models like TIMERS Mathur et al. (2021) and ConTempo Niu et al. (2024) adopt structured reasoning layers for multi-hop and document-level inference, often integrating temporal graphs or symbolic abstractions to guide event understanding.
To evaluate these capabilities more rigorously, new benchmarks have emerged. TRAM Wang and Zhao (2024) assesses LLMs on a spectrum of temporal tasks, including event frequency, duration estimation, and timeline ordering. Even state-of-the-art models like GPT-4 underperform on these tasks compared to human baselines, highlighting persistent gaps. Synthetic evaluations such as Test of Time (ToT) Fatemi et al. (2024) isolate core temporal reasoning skills like logic, arithmetic, and counterfactuals, free from the confounding effects of memorization. TODAY Feng et al. (2023) and Narrative-of-Thought Zhang et al. (2024) further probe a model’s ability to reason over temporal shifts and structured narratives.
Beyond accuracy, robustness remains a critical challenge. Studies such as Wallat et al. (2024, 2025) expose temporal blind spots in LLMs when applied on adversarial datasets, showing how small changes in timestamps or event order can significantly degrade performance, revealing a lack of temporal generalization and an over-reliance on surface form cues.
While temporal reasoning in PLMs has improved through targeted objectives and structured modeling, current systems still fall short in handling implicit cues, abstract reasoning, and robustness across time. Future work must explore deeper symbolic integration, stronger generalization under temporal drift, and evaluation setups that disentangle reasoning from retrieval and memorization.
Summary: TQA has progressed from rule-based pipelines to neural architectures that integrate temporal information through input representations, model internals, and dynamic retrieval.
TLMs improved contextualization by embedding timestamps and temporal cues, yet they remain constrained by static training corpora and a focus on explicit time expressions often falling short on vague, implicit, or underspecified temporal signals. Temporal RAG systems address some of these limitations by retrieving external, time-sensitive evidence at inference time, but their effectiveness depends heavily on the quality, relevance, and freshness of the retrieved content.
Recent reasoning-focused architectures go a step further, enabling event ordering and multi-hop inference, yet even state-of-the-art models exhibit temporal blind spots and struggle with robust generalization across temporal contexts.
Collectively, while each paradigm addresses specific limitations of its predecessor, none fully resolves the challenges of temporal ambiguity, drift, and reasoning. Future work must systematically target these failure modes, particularly the balance between memorized and retrieved knowledge, the handling of underspecified timeframes, and evaluation methods that isolate genuine temporal reasoning from pattern matching. Bridging these gaps is essential for building truly robust and temporally aware QA systems.
5 Future Research Directions
Despite significant advances in neural temporal reasoning, current TQA systems face fundamental limitations that require targeted research efforts. We identify the following critical areas where future work is needed to build temporally aware, trustworthy QA systems. Detailed technical challenges and potential solutions for each direction are provided in Appendix F.
Dynamic Temporal Knowledge Management.
Current TQA systems rely on static corpora, making them unable to respond to fast-evolving information needs. The temporal propagation problem, where small updates disrupt temporal dependencies across related events, exposes fundamental architectural limitations. Future systems must move beyond isolated fact updates toward scalable frameworks that track, edit, and reason over temporal dependencies in real time.
Temporally-Aware LLM Agents.
While LLMs demonstrate impressive general reasoning capabilities, they consistently underperform in temporal reasoning tasks. LLMs exhibit temporal hallucinations and fail to resolve context-dependent expressions such as "last Tuesday" or "since our previous discussion." Incorporating timeline tracking, event memory, and temporal reference resolution is essential for temporally coherent dialogue agents.
Diachronic and Synchronic Knowledge Integration.
Temporal questions often require combining time-sensitive sources showing change over time with stable sources summarizing information at specific points. Most current systems treat these separately, limiting their ability to answer questions involving both historical trends and current facts. Future systems should develop temporal alignment algorithms and cross-source reasoning frameworks.
Temporal Uncertainty and Confidence Modeling.
Many historical events have unclear start and end dates, yet most TQA systems treat all dates as exact. This mismatch creates problems when systems give confident answers about uncertain information, with issues compounding in multi-step reasoning. Future systems should handle uncertain dates and provide confidence scores for temporal answers.
Multilingual and Multimodal Temporal QA.
Temporal expressions, date formats, and cultural references vary widely across languages and modalities, yet most TQA systems are developed primarily for English textual input. Current systems struggle with diverse temporal signals due to limited cultural grounding and inadequate multimodal integration. Future research should focus on developing multilingual temporal taggers and cross-modal alignment techniques.
Implicit Temporal Intent Understanding.
Many temporal questions hide their time requirements rather than stating them directly. Current systems struggle with this ambiguity because they lack mechanisms to infer intended timeframes from contextual clues or shared knowledge. As users increasingly expect models to infer nuanced, context-sensitive answers, detecting hidden temporal assumptions becomes critical for avoiding plausible but misleading outputs.
Evaluation and Benchmarking for Temporal Reasoning.
Standard metrics like accuracy and F1 fall short in capturing temporal coherence or reasoning depth. Developing temporally-aware evaluation metrics remains an important open challenge for assessing and comparing future systems more effectively.
@@ 6 Conclusion
TQA plays a critical role in retrieving and reasoning over time-sensitive information in dynamic contexts. In this survey, we focus on recent advancements in the field, particularly modern TLMs and RAG approaches, examining core challenges including temporal expression recognition, event ordering, and implicit temporal reasoning.
Despite advances, current systems struggle with answer drift, temporal uncertainty, cultural and linguistic diversity, and limited temporal granularity. Most benchmarks rely on static snapshots, hindering real-time and future-oriented reasoning. Addressing these gaps requires better temporal representations, dynamic evaluation, adaptive learning.
@@
This survey aims to provide a comprehensive overview of TQA. There are a few important limitations to acknowledge.
We made our best efforts to be thorough, but it is possible that some relevant works may have been missed. We conducted an extensive literature review using forward and backward snowballing techniques, with particular attention to papers published in major venues such as ACL, SIGIR, EMNLP, NeurIPS, ECIR, and preprints on arXiv. On the other hand, due to page limitations, we provide only a brief summary of each method, without providing exhaustive technical details.
Appendix A Related Surveys
@@
While QA have been widely surveyed, most existing reviews focus on general techniques, often neglecting temporal aspects. IR surveys emphasize ranking functions, neural retrieval models, and query understanding Li et al. (2025); Zhu et al. (2023), while QA surveys center on extractive, abstractive, or multi-hop answering over static knowledge sources Zhu et al. (2021); Mavi et al. (2024). They rarely consider temporal intent, dynamic or evolving information, or event sequencing, highlighting a key gap that remains unaddressed.
Several earlier works provided foundational insights into TIR. Alonso et al. (2011) discusses challenges such as real-time streams, exploratory temporal search, and spatio-temporal retrieval. Campos et al. (2014) offers a broad overview of document dating, time-aware ranking, and query understanding, covering both explicit and implicit time signals. Kanhabua and Anand (2016) complements these with a tutorial on temporal indexing and ranking, emphasizing the detection of temporal query intent. There has been however no recent systematic overview of the field despite much research interest.
As a parallel research line, TQA over knowledge graphs has gained considerable attention Jia et al. (2021); Saxena et al. (2021); Chen et al. (2023b); Xiong et al. (2024b). Our survey focuses on temporally aware QA over text. We review both traditional and neural approaches to core tasks such as temporal tagging, event dating, time-aware retrieval, and temporal reasoning. To our knowledge, no prior survey brings together recent developments across these tasks in the context of text-based QA. Other related topics, including temporal fact verification Barik et al. (2024) and timeline summarization Sojitra et al. (2024), are discussed only when directly relevant.
@@
Temporal QA, draw=black, fill=gray!10, thick, text=black, ver
@@
Temporal Tasks ( §C),
@@
Das et al. (2017), Wang et al. (2021b) , Morbidoni et al. (2018), Wang et al. (2021b)
, color=cyan!100, fill=cyan!15, thick, text=black, text width=20em, leaf
@@
Dalli (2006), Jatowt et al. (2007), Kumar et al. (2012)
, color=cyan!100, fill=cyan!15, thick, text=black, text width=20em, leaf
@@
Jatowt et al. (2013), Shrivastava et al. (2017)
, color=cyan!100, fill=cyan!15, thick, text=black, text width=20em, leaf
@@
Kanhabua and Nørvåg (2010), Dakka et al. (2008), Jones and Diaz (2007)
, color=cyan!100, fill=cyan!15, thick, text=black, text width=20em, leaf